The waiting time X to get on a certain network can be 1, 5, 10 or 12 seconds. Find P(X = 5), P(X ≥ 2), P(3 < X < 12) if the probability mass function of X is:

x: 1 5 10 12

f(x): .6 .2 .1 .1 (respectively)
Answer #1

From given information we have

P(X=5) = 0.20

P(X>=2) = P(X=5) + P(X=10)+P(X=12)=0.2+0.1+0.1 = 0.40

P(3<X<12) = P(X=5) + P(X=10) = 0.2+0.1 = 0.30